### v3

# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Theory Perspective

Abstract
The Rectified Power Unit (RePU) activation functions, unlike the Rectified Linear Unit (ReLU), have the advantage of being a differentiable function when constructing neural networks. However, it can be experimentally observed when deep layers are stacked, neural networks constructed with RePU encounter critical issues. These issues include the values exploding or vanishing and failure of training. And these happen regardless of the hyperparameter initialization. From the perspective of effective theory, we aim to identify the causes of this phenomenon and propose a new activation function that retains the advantages of RePU while overcoming its drawbacks.
keywords:
Neural Networks , Activation Functions , Deep Learning Theory , Effective Theory , Rectified Power Units
journal: Neural Networks
[…]
organization=Department of Mathematical Sciences, Seoul National University,addressline=1, Gwanak-ro, Gwanak-gu, city=Seoul, postcode=08826, state=Seoul, country=South Korea
1 Introduction
1.1 Activation Functions
An activation function is a function that acts on the output of each layer of a neural network. The remarkable success of deep neural networks is closely related to the choice of appropriate nonlinear activation functions, and there has been extensive discussion solely on the research of activation functions. Early studies on neural networks include the perceptron [1], where the activation function is the sign function, and it was composed of only one or two layers. The perceptron had the drawback that its derivative was almost everywhere zero, making it difficult to apply the backpropagation algorithm. According to Cybenko [2], for a neural network to satisfy the universal approximation property, the activation function must be bounded, non-constant, monotonically increasing, and continuous. In the 1980s and 1990s, smooth activations like the sigmoid activation and tanh activation, along with their variations, were proposed [3], [4], [19]. These activations had the disadvantage of gradient vanishing when stacking layers deeply. To overcome this drawback, the Rectified Linear Unit (ReLU) was proposed by Nair and Hinton [5], whose derivative is a non-zero constant. Variations of ReLU, such as PReLU, GELU, and ELU, have also been proposed in the 2010s [7], [8], [9]. Recently, there have been attempts to find activations tailored to specific tasks to impart inductive bias to the neural network structure according to the task [10], [11].
1.2 Rectified Power Unit
One of the attempts to provide good regularity and inductive bias to the architecture is the Rectified Power Unit (RePU), a generalization of ReLU [12]. Early results using RePU include the application of ReQU, a type of RePU, in the DeepRitz method, which is a neural network approach for solving Partial Differential Equations (PDEs) [13]. RePU has also been proposed to construct differentiable neural networks, and error bounds related to this have been analyzed [12]. According to research, using a RePU network allows for the optimal approximation of smooth functions in Sobolev space with optimal depth and width, and the optimal approximation error bound is also provided in [12]. Similarly, another study derived that a shallow RePU network can approximate multivariate functions of a certain function class with finite neural network weight norms and possibly unbounded width [14]. Furthermore, it was shown that the derivatives of a network composed of RePU can be expressed as a mixture of various types of RePU, and an improved upper bound on the complexity of these derivatives was derived. This demonstrates the capability of RePU by showing the risk bound for the deep score matching estimator (DSME) [15]. Despite these results highlighting the good regularity of RePU neural networks, experimental evidence shows that RePU networks suffer from exploding or vanishing values and fail to train properly when the layers are deep.
1.3 Effective Theory of Neural Networks
The language of quantum field theory (QFT) is useful for analyzing systems with a very high degree of freedom. One such example is artificial neural networks. It is known that the distribution of neural network ensembles follows a Gaussian process when the width is infinite, corresponding to the free field in the QFT framework [16], [17], [18], [24], [25]. In practice, neural networks with finite width exhibit non-Gaussianity in their distributions as layers are stacked, allowing the distribution of neurons to be calculated perturbatively, as in weakly-interacting theory [24], [25]. Interestingly, when the width of the neural network is finite, the propagation of information through the neural network can be understood as a renormalization group (RG) flow. By analyzing the recursive relations of the kernels in RG flow, hyperparameters can be tuned to a criticality state that ensures the stability of the network [18]. Additionally, this analysis allows for the classification of activation functions into universality classes based on the behavior of the RG flow [18]. Some types of activation functions, according to this classification, cannot satisfy criticality conditions from the outset.
1.4 Our Contribution
We theoretically analyze the RePU activation from the perspective of effective theory. Specifically, we calculate the susceptibility of the RePU activation to theoretically demonstrate that RePU cannot satisfy the criticality condition. We then verify this experimentally by calculating the empirical kernel of the initialized neural network and during the training process. Not only do we identify the fundamental failure of RePU, but we also propose a modified RePU (MRePU) based on theoretical predictions to address these issues. We demonstrate the improvements experimentally by evaluating the empirical kernel and testing performance of MRePU, as compared to RePU.
2 Preliminary
In this section, we define the neural network and activation functions that will be the subject of our analysis, focusing particularly on the Rectified Power Unit (RePU). Additionally, we explore the effective theory of neural networks, which will be the primary tool for our analysis.
2.1 Overview of Neural Networks and Activation Functions
We primarily deal with neural networks having a fully connected network (FCN) architecture in this paper. An FCN is composed of affine transformations with hyperparameters weights and biases at each layer, followed by a nonlinear activation function applied component-wise. Specifically, it is structured as defined below:
Definition 1 (Fully Connected Network (FCN)) A Fully Connected Network $f(x;\theta):\mathbb{R}^{n_{in}}\rightarrow\mathbb{R}^{n_{out}}$ is defined by the following recursive equations:
$\begin{split}z_{i}^{(1)}(x_{\alpha})&:=\sum_{j=1}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}+b_{i}^{(1)},\quad\text{for}\quad i=1,\ldots,n_{1},\\ z_{i}^{(l+1)}(x_{\alpha})&:=\sum_{j=1}^{n_{l}}W_{ij}^{(l+1)}\sigma\Big{(}z_{j}^{(l)}(x_{\alpha})\Big{)}+b_{i}^{(l+1)},\quad\text{for}\quad i=1,\ldots,n_{l+1};\\ &\quad l=1,\ldots,L-1.\end{split}$ (1)
where $n_{in}$ is the input dimension of the neural network, $n_{out}$ is the output dimension of the neural network, and each $n_{l}$ is the width (i.e., the number of nodes at the $l$-th layer). $L$ is the depth of stacked layers. The parameters $(b_{i}^{(l)})_{i=1,...,n_{l}}$, $(W_{ij}^{(l)})_{i=1,...,n_{l},j=1,...,n_{l-1}}$ are bias vectors and weight matrices, respectively. $z_{i}^{(l)}$ is called the preactivation at the $l$-th layer. $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activation function that acts on each node of the preactivations.
In the previous definition of neural networks, we also explained what an activation function is. The connections between the nodes of a neural network can be seen as a type of graph, and this graph representing the connections of the neural network is called the architecture. Besides the architecture, there are several factors that determine the properties of a neural network (training dynamics, inductive bias, approximation properties, etc.). In this paper, we focus on the initialization of the weight and bias parameters and the activation function.
Definition 2 (Initialization distribution of biases and weights) In the definition of neural networks, there are adjustable weights and biases parameters. The method of setting these parameters at the beginning of training is called initialization, and typically, each parameter follows a specific probability distribution, referred to as the initialization distribution. Although it is possible to use mathematically complex initialization distributions, it is common practice to assign each weight and bias under the condition of being independent and identically distributed (i.i.d). The probability distributions used can include Gaussian distribution, truncated Gaussian distribution, uniform distribution, and others. For wide neural networks, it is known that when the mean of the weight distribution is 0 and the variance is constant, differences arising from different types of distributions are suppressed by $\frac{1}{width}$. Therefore, for the sake of convenience in our discussion, we will set the distributions of weights and biases to follow a mean-zero Gaussian distribution:
$\begin{split}\mathbb{E}\Big{[}b_{i}^{(l)}b_{j}^{(l)}\Big{]}&=\delta_{ij}C_{b}^{(l)}\\ \mathbb{E}\Big{[}W_{i_{1}j_{1}}^{(l)}W_{i_{2}j_{2}}^{(l)}\Big{]}&=\delta_{i_{1}i_{2}}\delta_{j_{1}j_{2}}\frac{C_{W}^{(l)}}{n_{l-1}}.\end{split}$ (2)
where the set of bias variances $\{C_{b}^{(1)},\ldots,C_{b}^{(L)}\}$ and the set of rescaled weight variances $\{C_{W}^{(1)},\ldots,C_{W}^{(L)}\}$ are called initialization hyperparameters.
Activation functions Since the inception of artificial neural networks, various types of activation functions have been proposed. An activation function is essentially a 1D real scalar-valued function that acts on the preactivation, which is the output of a node. Because the capabilities of a neural network are greatly influenced by the type of activation function used, extensive research has been conducted in this area. We provide a brief summary of the activation functions proposed so far through Table 1.
Name Formula References
Perceptron $\sigma(z)=\begin{cases}1\text{,}&\quad\text{if }z\geq 0\\ 0\text{,}&\quad\text{if }z<0\\ \end{cases}$ [1]
Sigmoid $\sigma(z)=\frac{1}{1+e^{-z}}$ [2]
Tanh $\sigma(z)=\tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ [2], [19]
Sin $\sigma(z)=sin(z)$ [20], [21]
ReLU $\sigma(z)=\begin{cases}z\text{,}&\quad\text{if }z\geq 0\\ 0\text{,}&\quad\text{if }z<0\\ \end{cases}$ [5], [6]
Leaky ReLU $\sigma(z)=\begin{cases}z\text{,}&\quad\text{if }z\geq 0\\ \alpha z\text{,}&\quad\text{if }z<0\\ \end{cases}$ [7], [22]
Softplus $\sigma(z)=\log(1+e^{z})$ [23]
SWISH $\sigma(z)=\frac{z}{1+e^{-z}}$ [10]
GELU $\sigma(z)=\Big{[}\frac{1}{2}+\frac{1}{2}\text{erf}\Big{(}\frac{z}{\sqrt{2}}\Big{)}\Big{]}z$ [8]
Table 1: Various Kinds of Activation functions
Now we define the RePU activation functions, which are the main objective of our work.
Definition 3 (Rectified Power Unit (RePU) The RePU activation function is defined by the following equation:
$\begin{split}\sigma(z)=\begin{cases}z^{p}\text{,}&\quad\text{if }z\geq 0\\ 0\text{,}&\quad\text{if }z<0\\ \end{cases}.\end{split}$
where $p$ is a positive integer that determines the power to which the input $z$ is raised. When $p=1$, it corresponds to ReLU, and thus RePU can be considered a generalized version of ReLU. The cases where $p=2$ and $p=3$ are referred to as Rectified Quadratic Unit (ReQU) and Rectified Cubic Unit (ReCU), respectively. For RePU with $p=k$, it is easy to see that it is differentiable $k-1$ times, and this property ensures that the neural network and its gradients are differentiable functions.
2.2 Effective Field Theory for Neural Networks
We utilize the framework of effective field theory to understand the distribution of preactivations in a neural network ensemble sampled from the initialization distribution. To accomplish this, we first define the necessary concepts.
Definition 4 ($M$-point correlators) The moments or $M$-point correlators of a probability distribution $p(z)=p(z_{1},...,z_{N})$ are defined as follows:
[…]
The collection of all $M$-point correlators fully characterizes the probability distribution since the information from the M-point correlators allows us to compute the expected values of analytic observables. For a Gaussian distribution, it is important to note that the distribution’s information can be fully specified by the 1-point correlator (mean) and the 2-point correlator (variance). Keeping this in mind, we slightly modify the $M$-point correlator to define what is known as the connected correlator (also known as cumulant). The definition is as follows:
Definition 5 ($M$-point connected correlators) For $2\leq M$, the cumulant or $M$-point connected correlators of a probability distribution $p(z)$ are defined as follows:
$\begin{split}&\mathbb{E}[z_{\mu_{1}}\cdots z_{\mu_{M}}]|_{\text{connected}}=\mathbb{E}[z_{\mu_{1}}\cdots z_{\mu_{M}}]-\\ &\sum_{\text{all subdivisions of }(\mu_{1},...,\mu_{M})}\mathbb{E}[z_{\mu_{i_{1,1}}}\cdots z_{\mu_{i_{1,j_{1}}}}]|_{\text{connected}}\cdots\mathbb{E}[z_{\mu_{i_{k,1}}}\cdots z_{\mu_{i_{k,j_{k}}}}]|_{\text{connected}}.\end{split}$
For $M=1$, the connected correlator coincides with the regular correlator. For parity-symmetric distributions, it also coincides for $M=2$. According to Wick’s theorem, it is known that for a Gaussian distribution, the connected correlators for $M>2$ are zero. Therefore, higher-point connected correlators can be used to determine how much the distribution deviates from a Gaussian distribution. Additionally, if the connected correlators for $M>2$ are small, the distribution is defined as nearly-Gaussian. In quantum field theory, the 2-point connected correlator can be seen as an expression of the translational symmetry of a free field. In a free field, there is no interaction due to potential, so all connected correlators higher than the 2-point vanish. Conversely, if there are interactions between particles due to a potential, higher-order connected correlators become non-zero. Similarly, in the neural network we analyze, if the higher-order correlators do not vanish and instead show significant values, this implies meaningful interactions between nodes in the feature space, suggesting that feature learning is taking place.
Using these concepts, we can analyze the probability distribution of the preactivations of an ensemble following the initialization distribution. Specifically, since each layer depends on the previous layer, marginalization reveals that there are recursive relations between the statistics of the probability distributions of the layers.
Induced distributions Following the notation in (1), let $p(z^{(l)}|\mathcal{D})$ denote the probability distribution of the preactivations in the $l$-th layer given a dataset $\mathcal{D}=\{x_{i;\alpha}\}_{i=1,...,n_{0};\alpha=1,...,N_{\mathcal{D}}}$. For the first layer, the preactivation has $M>2$ connected correlators equal to 0, and has the following mean and covariance, indicating that each preactivation follows an independent Gaussian distribution.
$\begin{split}\mathbb{E}[z_{i;\alpha}^{(1)}]&=\mathbb{E}\bigg{[}b_{i}^{(1)}+\sum_{j=0}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}\bigg{]}=0,\\ \mathbb{E}[z_{i_{1};\alpha_{1}}^{(1)}z_{i_{2};\alpha_{2}}^{(1)}]&=\mathbb{E}\bigg{[}\bigg{(}b_{i_{1}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{1}j_{1}}^{(1)}x_{j_{1};\alpha_{1}}\bigg{)}\bigg{(}b_{i_{2}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{2}j_{2}}^{(1)}x_{j_{2};\alpha_{2}}\bigg{)}\bigg{]}\\ &=\delta_{i_{1}i_{2]}}\bigg{(}C_{b}^{(1)}+C_{W}^{(1)}\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}x_{j;\alpha_{1}}x_{j;\alpha_{2}}\bigg{)}=\delta_{i_{1}i_{2}}G_{\alpha_{1}\alpha_{2}}^{(1)}.\end{split}$
Where $G_{\alpha_{1}\alpha_{2}}^{(1)}$ is called as the metric, which represents the two-point correlator of the preactivations for different samples. Similarly, the conditional probability $p(z^{(l+1)}|z^{(l)})$ is also described by Gaussian distribution with covariance $\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$, where the distribution (i.e.$\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$) depends on the condition $z^{(l)}$ We refer to this $\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}$ as the $l+1$-th layer stochastic metric. We can consider the mean metric $G_{\alpha_{1}\alpha_{2}}^{(l+1)}$, which is the average of the stochastic metric. Additionally, we can consider the variance of $\Delta\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}=\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}-G_{\alpha_{1}\alpha_{2}}^{(l+1)}$, which is the fluctuation of the stochastic metric around the mean metric. We refer to scaled quantity of this variance as the four-point vertex, defined as follows:
$V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}:=n_{l}\mathbb{E}\bigg{[}\Delta\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}\Delta\hat{G}_{\alpha_{3}\alpha_{4}}^{(l+1)}\bigg{]}=n_{l}\bigg{(}\mathbb{E}\bigg{[}\hat{G}_{\alpha_{1}\alpha_{2}}^{(l+1)}\hat{G}_{\alpha_{3}\alpha_{4}}^{(l+1)}\bigg{]}-G_{\alpha_{1}\alpha_{2}}^{(l+1)}G_{\alpha_{3}\alpha_{4}}^{(l+1)}\bigg{)}.$
The four-point vertex is related to the 4-point connected correlator as follows:
$\begin{split}&\mathbb{E}[z_{i_{1};\alpha_{1}}^{(l+1)}z_{i_{2};\alpha_{2}}^{(l+1)}z_{i_{3};\alpha_{3}}^{(l+1)}z_{i_{4};\alpha_{4}}^{(l+1)}]|_{\text{connected}}\\ &=\frac{1}{n_{l}}[\delta_{i_{1}i_{2}}\delta_{i_{3}i_{4}}V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}+\delta_{i_{1}i_{3}}\delta_{i_{2}i_{4}}V^{(l+1)}_{(\alpha_{1}\alpha_{3})(\alpha_{2}\alpha_{4})}+\delta_{i_{1}i_{4}}\delta_{i_{2}i_{3}}V^{(l+1)}_{(\alpha_{1}\alpha_{4})(\alpha_{2}\alpha_{3})}].\end{split}$ (3)
The quantities we have just defined contain significant information about the distribution of the preactivation $p(z^{(l)}|\mathcal{D})$. If we define the action $S$ of $p(z^{(l)}|\mathcal{D})$ as follows (parity symmetry can be easily checked):
$S:=\frac{1}{2}\sum_{i=1}^{n_{l}}\sum_{\alpha_{1}\alpha_{2}\in\mathcal{D}}g^{\alpha_{1}\alpha_{2}}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}-\frac{1}{8}\sum_{i_{1},i_{2}=1}^{n_{l}}\sum_{\alpha_{1},...,\alpha_{4}\in\mathcal{D}}v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}z_{i;\alpha_{1}}^{(l)}z_{i;\alpha_{2}}^{(l)}z_{i;\alpha_{3}}^{(l)}z_{i;\alpha_{4}}^{(l)}+\cdots.$
It is known that the following relationship holds for the coefficients. This can be derived by comparing the coefficients of the connected correlator given by the mean metric and four-point vertex with those of the connected correlator assumed by the ansatz provided by action $S$. $g^{\alpha_{1}\alpha_{2}}_{(l+1)}$ and $v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l+1)}$ [18]:
$\begin{split}g^{\alpha_{1}\alpha_{2}}_{(l)}&=G^{\alpha_{1}\alpha_{2}}_{(l)}+O(\frac{1}{n_{l-1}}),\\ v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}&=\frac{1}{n_{l}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}+O(\frac{1}{n_{l-1}^{2}}).\end{split}$
Where, $G^{\alpha_{1}\alpha_{2}}_{(l)}$ is inverse of mean metric $G_{\alpha_{1}\alpha_{2}}^{(l)}$ and $V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}$ is defined as follows:
$V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}:=\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}G^{\alpha_{1}\beta_{1}}_{(l)}G^{\alpha_{2}\beta_{2}}_{(l)}G^{\alpha_{3}\beta_{3}}_{(l)}G^{\alpha_{4}\beta_{4}}_{(l)}V^{(l)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}.$
Kernels and Representation Group flow Now, we will investigate how the statistics of the preactivation change as they pass through the layers. We will refer to the evolution of the distribution through the layers as the representation group flow, following the terminology of [18]. While it is not feasible to explicitly calculate the statistics for a complex non-Gaussian distribution, it is possible to develop a perturbative theory to understand the recursive flow of the key statistics of the distribution. Before delving into this, let’s define the notation for Gaussian expectations. For an observable $\mathcal{O}(z)$ with a mean-zero Gaussian distribution with variance $\bf{K}=(K_{ij})$, the expectation is denoted as $\langle\mathcal{O}(z)\rangle_{\bf{K}}$. Now, if we define the kernel of the neural network through the following recursive relation:
$\begin{split}K^{(1)}_{\alpha_{1}\alpha_{2}}&:=G^{(1)}_{\alpha_{1}\alpha_{2}},\\ K^{(l+1)}_{\alpha_{1}\alpha_{2}}&=C^{(l+1)}_{b}+C^{(l+1)}_{W}\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}.\end{split}$
it is known that this kernel satisfies following properties, as discussed in [18].
$\begin{split}G^{(l)}_{\alpha_{1}\alpha_{2}}&=K^{(l)}_{\alpha_{1}\alpha_{2}}+O(\frac{1}{n_{l-1}}),\\ V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}&=\Big{(}C_{W}^{(l+1)}\Big{)}^{2}\Big{[}\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &-\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}-\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &+\frac{n_{l}}{4n_{l-1}}\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}\Big{\langle}\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\Big{(}z_{\beta_{1}}z_{\beta_{2}}-K_{\beta_{1}\beta_{2}}^{(l)}\Big{)}\Big{\rangle}_{K^{(l)}}\\ &\times\Big{\langle}\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\Big{(}z_{\beta_{3}}z_{\beta_{4}}-K_{\beta_{3}\beta_{4}}^{(l)}\Big{)}\Big{\rangle}_{K^{(l)}}\Big{]}+O(\frac{1}{n_{l}}).\end{split}$ (4)
Susceptibility When expanding the kernel matrix for two inputs $(x_{i;+})$ and $(x_{i;-})$, it can be written as follows:
$K^{(l)}=\begin{pmatrix}K^{l}_{++}&K^{l}_{+-}\\ K^{l}_{-+}&K^{l}_{--}\end{pmatrix}=K^{(l)}_{[0]}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}+K^{(l)}_{[1]}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}+K^{(l)}_{[2]}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}.$
Where, $K^{(l)}_{\pm\pm}:=\mathbb{E}\bigg{[}\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}(z_{i;\pm}^{(l)})^{2}\bigg{]}$, $K^{(l)}_{\pm\mp}:=\mathbb{E}\bigg{[}\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}z_{i;\pm}^{(l)}z_{i;\mp}^{(l)}\bigg{]}$ The coefficients $K^{(l)}_{[\alpha]}$ are given by:
$\begin{split}K_{[0]}^{(l)}&=\frac{1}{4}[K^{(l)}_{++}+K^{(l)}_{--}+2K^{(l)}_{+-}],\\ K_{[1]}^{(l)}&=\frac{1}{2}[K^{(l)}_{++}-K^{(l)}_{--}],\\ K_{[2]}^{(l)}&=\frac{1}{4}[K^{(l)}_{++}+K^{(l)}_{--}-2K^{(l)}_{+-}].\end{split}$
If we express $(x_{i;+})$ and $(x_{i;-})$ as $x_{i;\pm}=\frac{x_{i;+}+x_{i;-}}{2}\pm\delta x_{i}$, it can be seen that as $(\delta x_{i})$ approaches zero, the coefficients $K^{(l)}_{[1]},K^{(l)}_{[2]}$ also approach zero. Consequently, the coefficients can be expanded as follows:
$\begin{split}K_{[0]}^{(l)}&=K_{00}^{(l)}+\delta^{2}K_{[0]}^{(l)}+O(\delta^{4}),\\ K_{[1]}^{(l)}&=\delta K_{[1]}^{(l)}+\delta^{3}K_{[1]}^{(l)}+O(\delta^{5}),\\ K_{[2]}^{(l)}&=\delta^{2}K_{[2]}^{(l)}+\delta^{4}K_{[2]}^{(l)}+O(\delta^{6}).\end{split}$
Then, it is known that the following recursion can be derived [18]:
$\begin{split}K_{00}^{(l+1)}&=C_{b}+C_{W}g(K_{00}^{(l)}),\\ \delta K_{[1]}^{(l+1)}&=\mathcal{X}_{\parallel}(K_{00}^{(l)})\delta K_{[1]}^{(l)},\\ \delta^{2}K_{[2]}^{(l+1)}&=\mathcal{X}_{\perp}(K_{00}^{(l)})\delta^{2}K^{(l)}_{[2]}+h(K^{(l)}_{00})(\delta K^{(l)}_{[1]})^{2}.\end{split}$
Where $K_{00}^{(l+1)}$ is the kernel for a single input, and each $g$, $h$, $\mathcal{X}_{\parallel}$, $\mathcal{X}_{\perp}$ are defined as follows:
$\begin{split}g(K)&:=\langle\sigma(z)\sigma(z)\rangle_{K},\\ \mathcal{X}_{\parallel}(K)&:=C_{W}g^{\prime}(K)=\frac{C_{W}}{K}\langle z\sigma^{\prime}(z)\sigma(z)\rangle_{K},\\ \mathcal{X}_{\perp}(K)&:=C_{W}\langle\sigma^{\prime}(z)\sigma^{\prime}(z)\rangle_{K},\\ h(K)&:=\frac{C_{W}}{4K^{2}}\Big{\langle}\sigma^{\prime}(z)\sigma^{\prime}(z)\Big{(}z^{2}-K\Big{)}\Big{\rangle}_{K}=\frac{1}{2}\frac{d}{dK}\mathcal{X}_{\perp}(K).\end{split}$ (5)
$K_{[1]}^{(l)}$ and $K_{[2]}^{(l)}$ represent the difference in the squared distances between two inputs and a value related to the square of the difference between the two inputs, respectively. As the inputs become closer, their leading terms exhibit a geometric ratio in susceptibility as they pass through layers, which forms the basis for naming $\mathcal{X}_{\parallel}$ and $\mathcal{X}_{\perp}$ as the parallel and perpendicular susceptibility, respectively. Since each leading term follows this geometric ratio in susceptibility, if the value is less than 1, the kernel vanishes along the layer, and if it is greater than 1, it diverges. Thus, setting each susceptibility to 1 is essential for maintaining a stable kernel.
3 Failure of RePU Activation
3.1 Susceptibility Calculation
Activation functions that exhibit similar behaviors according to the RG flow are classified into universality classes. Universality class is a concept in statistical physics used to classify systems that exhibit similar behavior at the limit scale. According to [18], scale-invariant activations (e.g. ReLU, LeakyReLU) form a single universality class called the self-invariant universality class. Tanh and Sin fall into the $K^{\star}=0$ universality class, while SWISH and GELU belong to the half-stable universality class. On the other hand, Perceptron, Sigmoid, and Softplus are known to not satisfy criticality at all. We will now calculate the susceptibility for the RePU activation function following equation (5) and demonstrate that RePU also fails to satisfy criticality. For the RePU activation function of order $p$, the parallel susceptibility and perpendicular susceptibility are calculated as follows:
[…]
The parallel susceptibility to perpendicular susceptibility ratio is $2p-1:p$. Therefore, except for $p=1$ (i.e., ReLU), it is impossible to set both susceptibilities to 1 simultaneously. If we set the parallel susceptibility to 1, the perpendicular susceptibility will be less than 1, causing the distance between two inputs to vanish as they propagate through the layers. Conversely, if we set the perpendicular susceptibility to 1, the parallel susceptibility will be greater than 1, causing each input to diverge as it propagates through the layers.
3.2 Experimental Validation
Empirical Kernel Behavior at Initialization According to the results in section 3.1, the kernel value of Neural Networks with RePU (for $p>1$) for a single input must either explode or vanish as it propagates through the layers under the initialization distribution. In this section, we experimentally verify this. For the experiment, the neural network is configured with an input dimension of 2, an output dimension of 1, hidden layers with widths of 512, and a depth of 10 (therefore, the total number of layers, including the input and output layers, is 12). The activation function is RePU. The initialization distribution follows equation (2), with $C_{b}$ set to 0, and $C_{W}$ adjusted during the experiment. To experimentally observe the change in the kernel, we define the following empirical kernel quantities:
$\begin{split}\hat{K}^{(l)}(\Theta)|_{\mathcal{D}=x_{0}}:=\frac{1}{n_{l}}\sum_{i}^{n_{l}}\Big{(}z^{(l)}_{i}(\Theta)|_{\mathcal{D}=x_{0}}\Big{)}^{2}\\ \hat{K}^{(l)}(\mathcal{D})|_{\Theta=\Theta_{0}}:=\frac{1}{n_{l}}\sum_{i}^{n_{l}}\Big{(}z^{(l)}_{i}(\mathcal{D})|_{\Theta=\Theta_{0}}\Big{)}^{2}.\end{split}$
Where $z^{(l)}_{i}(\Theta)|_{\mathcal{D}=x_{0}}$ is a random variable conditioned on the input $x_{0}$ and has randomness on $\Theta$. Similarly, $z^{(l)}_{i}(\mathcal{D})|_{\Theta=\Theta_{0}}$ is a random variable conditioned on the parameter $\Theta$ and has randomness on the data input. Each component of sample in $\mathcal{D}$ follows $N(0,1)$. $x_{0}$ is also generated from $\mathcal{D}$. The experimental results for a width of 512, $p=2$, and $C_{W}=21.3$ are shown in Figures 1. We observed that if $C_{W}$ is set smaller, the kernel vanishes on an $\mathcal{O}(\exp(\exp(-l)))$ scale, and if $C_{W}$ is set larger, it explodes on an $\mathcal{O}(\exp(\exp(l)))$ scale. The results in the figure for $C_{W}=21.3$ show that the kernel is highly sensitive to initial conditions and tends to either double exponentially explode or vanish. According to equations (5), (6) the susceptibility itself acts as a geometric ratio and is proportional to the kernel. As this ratio increases exponentially, it causes the kernel to grow double exponentially.
Figure 1: Empirical Kernels at Initialization Across Layers for RePU Activation with $p=2$. Left: Data is fixed and randomness is in the weight parameters. Right: Weight parameters are fixed and data is random. Each line represents a sample.
Dynamics of Mean Outputs and Kernels during Training In this section, we experimentally observe the dynamics of kernels in RePU neural networks during training. In the backpropagation algorithm used for updating network parameters, each factor in the chain rule is proportional to the perpendicular susceptibility at each layer, as described by the following equation.
$\begin{split}\mathbb{E}\bigg{[}\sum_{j_{1},j_{2}}\frac{dz_{i}^{(l+1)}}{dz_{j_{1}}^{(l)}}\frac{dz_{i}^{(l+1)}}{dz_{j_{2}}^{(l)}}\bigg{]}&=\mathbb{E}\bigg{[}\sum_{j_{1},j_{2}}W_{ij_{1}}^{(l+1)}W_{ij_{2}}^{(l+1)}{\sigma^{\prime}}_{j_{1}}^{(l)}{\sigma^{\prime}}_{j_{2}}^{(l)}\bigg{]}\\ &=C_{W}\langle\sigma^{\prime}(z)\sigma^{\prime}(z)\rangle_{K^{(l)}}=C_{W}\mathcal{X}_{\perp}^{(l)}\end{split}$
Through this, we can infer that in RePU neural networks, the kernel either increases or decreases double exponentially, and since the perpendicular susceptibility is proportional to this, the training dynamics become highly unstable or converge rapidly. To experimentally verify this, we designed an experiment using synthetic data. The task involved observing how well the neural network performs regression on synthetic data, where the input data are points in the plane $(x,y)$ with $x\sim N(0,1)$ and $y\sim N(0,1)$, and the target is set to $x^{2}+y^{2}$. The training dataset is composed of 100 samples. We conducted experiments with a network width of 512 and with 1, 3, and 5 hidden layers, testing cases where $p=2$ and $p=3$. The $C_{W}$ values were adjusted based on the architecture; if $C_{W}$ is too small, the kernel decreases double exponentially, and if $C_{W}$ is too large, the kernel becomes highly unstable and eventually diverges. Therefore, we experimentally determined and set an appropriate $C_{W}$ value. We calculated the kernel for an ensemble of 100 neural networks, obtaining the mean and variance while excluding models where the kernel overflowed due to explosion. The results are presented in Tables 2 and 3, with the results for 5 hidden layers and $p=2$ depicted in Figure 2. We then tested the ensemble of 100 neural network models trained under each setting on a test dataset generated from the same distribution as the training data. The results are depicted in Figure 3. From the aggregated results of the kernel and the mean outputs on the test dataset, we observed that significant learning only occurs when there is a single hidden layer. When the network has 3 or 5 hidden layers, the kernel values quickly diminish as the layers deepen, leading to poor learning performance. Additionally, as the depth of the network increases, the mean output converges to a constant, with its standard deviation also converging to a very small value.
Epochs $(N_{h},C_{W})$ $(1,2.5)$ $(3,3.0)$ $(5,2.6)$ At Initialization $(1.39\text{*}10^{-3},3\text{*}10^{-4})$ $(1.28\text{*}10^{-6},1.5\text{*}10^{-6})$ $(6.91\text{*}10^{-11},1.9\text{*}10^{-10})$ At 250 Epoch $(1.71\text{*}10^{-3},4\text{*}10^{-4})$ $(2.80\text{*}10^{-5},1.2\text{*}10^{-5})$ $(4.29\text{*}10^{-7},2,7\text{*}10^{-7})$ At 500 Epoch $(2.71\text{*}10^{-3},8\text{*}10^{-4})$ $(7.44\text{*}10^{-5},5.3\text{*}10^{-5})$ $(1.03\text{*}10^{-6},7.2\text{*}10^{-7})$ At 750 Epoch $(4.57\text{*}10^{-3},1.5\text{*}10^{-3})$ $(1.35\text{*}10^{-4},9.7\text{*}10^{-5})$ $(1.75\text{*}10^{-6},1.3\text{*}10^{-6})$ At 1000 Epoch $(6.70\text{*}10^{-3},1.6\text{*}10^{-3})$ $(2.14\text{*}10^{-4},1.4\text{*}10^{-4})$ $(2.62\text{*}10^{-6},2.1\text{*}10^{-6})$
Table 2: Means and Standard Deviations of the Empirical Kernel at the End of Hidden Layers. Here, $N_{h}$ denotes the number of hidden layers, and $C_{W}$ represents hyperparameter for the variance of weight parameters at initialization. The order of RePU is 2 in this case.
Epochs $(N_{h},C_{W})$ $(1,4.5)$ $(3,2.9)$ $(5,2.7)$ At Initialization $(9.01\text{*}10^{-3},4\text{*}10^{-3})$ $(3.49\text{*}10^{-10},1.3\text{*}10^{-9})$ $(1.22\text{*}10^{-14},8.3\text{*}10^{-14})$ At 250 Epoch $(1.05\text{*}10^{-2},4\text{*}10^{-3})$ $(1.34\text{*}10^{-7},1.6\text{*}10^{-7})$ $(1.54\text{*}10^{-12},9.0\text{*}10^{-12})$ At 500 Epoch $(1.14\text{*}10^{-2},5.1\text{*}10^{-3})$ $(2.30\text{*}10^{-7},2.2\text{*}10^{-7})$ $(1.91\text{*}10^{-12},1.0\text{*}10^{-11})$ At 750 Epoch $(1.24\text{*}10^{-2},6.1\text{*}10^{-3})$ $(3.31\text{*}10^{-7},2.9\text{*}10^{-7})$ $(2.35\text{*}10^{-12},1.2\text{*}10^{-11})$ At 1000 Epoch $(1.36\text{*}10^{-2},7.5\text{*}10^{-3})$ $(4.53\text{*}10^{-7},4.0\text{*}10^{-7})$ $(2.81\text{*}10^{-12},1.5\text{*}10^{-11})$
Table 3: Means and Standard Deviations of the Empirical Kernel at the End of Hidden Layers. Here, $N_{h}$ denotes the number of hidden layers, and $C_{W}$ represents hyperparameter for the variance of weight parameters at initialization. The order of RePU is 3 in this case.
Figure 2: The evolution of the mean of empirical kernels over an ensemble of 100 models for $x_{0}=(1,0)$ as training progresses for the RePU activation with $p=2$. The shaded areas represent the region between $\log_{10}(\text{mean}\pm 0.1\times\text{standard deviation})$.
Figure 3: Mean outputs over an ensemble of 100 models versus target values on random test data. The shaded areas represent the region of 1 standard deviation. Top left: hidden layers ($N_{h}$)= 1, Top right: $N_{h}$ = 3, Bottom: $N_{h}$ = 5.
4 Modified Rectified Power Unit (MRePU)
In this section, we propose the modified Rectified Power Unit (MRePU) function to overcome the limitations of RePU, which cannot be stacked deeply. MRePU retains the advantages of RePU (e.g., differentiability) while fundamentally addressing the existing issues. The definition of MRePU is as follows:
Definition 6 (Modified Rectified Power Unit (MRePU) The MRePU activation function of order $p>1$ is defined by the following equation:
$\begin{split}\sigma_{m;p}(z)=\begin{cases}z(z+1)^{p}\text{,}&\quad\text{if }z\geq-1\\ 0\text{,}&\quad\text{if }z<-1\\ \end{cases}.\end{split}$
Similar to RePU, the MRePU function of order $p$ has the property of being differentiable $p-1$ times.
4.1 Numerical Calculation of Susceptibility
Here, we calculate the parallel and perpendicular susceptibilities of the MRePU function we proposed. The susceptibility of the MRePU function of order p can be expressed by the following equation:
[…]
Since getting value of (7) analytically is impossible, we calculate the susceptibility numerically. We fixed the value of $C_{W}$ at $1.0$ and plotted graphs by exponentially increasing the value of $K$ for cases where $p=2$ and $p=3$. The results can be seen in Figure 4, where it is observed that the ratio of the two susceptibilities converges to 1 as the value of $K$ approaches 0. This confirms that the MRePU activation belongs to the $K^{\star}=0$ universality class, and by appropriately adjusting the $C_{W}$ value, it is possible to ensure that the kernel neither explodes nor vanishes.
Figure 4: Parallel and Perpendicular Susceptibility and Their Ratio for MRePU with $p=2$ (Left) and $p=3$ (Right) with $(C_{W},C_{b})=(1,0)$
4.2 Experimental Results
Now, as we did with RePU Neural Networks, we will investigate the statistical properties of MRePU under the same experimental settings.
Empirical Kernel and Four-point Vertex Behavior at Initialization As with the RePU case, we investigate the statistical properties of the neural network ensemble given by the initialization distribution. The experimental environment remains identical to that of the RePU Networks, with the only change being the activation function, which is now MRePU with $p=2$. The results for $C_{W}=1.0$ are shown in Figure 5. As can be seen from the graph, the empirical kernel maintains a consistent scale across both random data and weight parameters. Additionally, we also examined the behavior of the Four-point vertex for MRePU. Equation (4) can be simplified for a single input as follows:
$V^{(l+1)}=\mathcal{X}_{\parallel}^{2}(K^{(l)})V^{(l)}+C_{W}^{2}\Big{[}\langle\sigma^{4}(z)\rangle_{K^{(l)}}-\langle\sigma^{2}(z)\rangle_{K^{(l)}}^{2}\Big{]}+O(\frac{1}{n_{l}}).$ (8)
Figure 6 illustrates how the empirical four-point vertex, calculated from an ensemble of 1000 models using equation (3), changes across layers (with input $x_{0}=(1,0)$). $\mathcal{X}_{\parallel}$ increases as the kernel grows, and since the kernel itself depends on $C_{W}$, the inductive change in the four-point vertex is determined by $C_{W}$ according to equation (8). As evidence of this, the left figure shows that the four-point vertex increases more steeply as $C_{W}$ becomes larger. We also investigated the behavior of the four-point vertex across different widths. The results indicate that with smaller widths, perturbative effects in equation (8) become more pronounced, leading to a more rapid increase in the four-point vertex, and overall, higher values are observed. The four-point vertex serves as an indicator of non-Gaussianity, which is expected to increase as more layers with non-linear activations are added. Moreover, since smaller widths imply stronger interactions between neurons, these experimental results align well with our theoretical predictions.
Figure 5: Empirical Kernels at Initialization Across Layers for MRePU Activation with $p=2$. Left: Data is fixed and randomness is in the weight parameters. Right: Weight parameters are fixed and data is random. Each line represents a sample.
Figure 6: Empirical Four-Point Vertex Across Layers at Initialization. Each value is calculated over an ensemble of 1000 models. Left: Width is fixed at 512, and $C_{W}$ is varying. Right: $C_{W}$ is fixed at $1.0$, and the width is varying.
Dynamics of Statistics During Training
Now, we will examine how the statistics of the MRePU Network ensemble change under training dynamics. The experimental setup is the same as in the case of the RePU Network, and the number of hidden layers is set to 5. In this case, we constructed the training dataset in the same way as in the RePU activation case, but the number of samples is set to 1000. Although the theoretical approximation suggests setting $C_{W}=1.0$, meaningful learning was only achieved with a slightly lower value, possibly due to the perturbative effects in the MRePU Network. In our experiments, we set $C_{W}=0.8$. The dynamics of the kernel and four-point vertex are depicted in Figure 7, while the comparison of the ensemble’s mean output against the target for the test dataset is shown in Figure 8. Surprisingly, we observed that even with a vanilla FCN (Fully Connected Network) without considering architectures like ResNet, which are designed to prevent issues like the vanishing gradient, the deep neural network with 5 hidden layers still achieved meaningful learning using the MRePU function we designed. Additionally, as shown on the left side of Figure  7, we also calculated the perpendicular susceptibility, which showed a tendency to approach 1 as training progressed. On the right side of the same figure, it can be seen that the ensemble’s four-point vertex decreases to 0 during initialization but tends to increase linearly with the layers as training progresses. An increasing four-point vertex indicates growing interactions between neurons, which implies that representation learning is occurring. This suggests that the deep vanilla MRePU network is indeed engaging in meaningful representation learning.
Figure 7: The evolution of the mean of empirical kernels (solid line, left y-axis) and perpendicular susceptibility (dotted line, right y-axis) (Left) and the four-point vertex (Right) over an ensemble of 100 models for $x_{0}=(1,0)$ as training progresses for the MRePU activation with p=2 . The shaded areas represent the region between $\log_{10}(\text{mean}\pm\text{standard deviation})$ (for Left).
Figure 8: Mean outputs over an ensemble of 100 models versus target values on random test data. The shaded areas represent the region of 1 standard deviation. The number of hidden layers is 5.
5 Approximation Properties of MRePU networks
In this section, we theoretically demonstrate that a neural network constructed with the MRePU activation function can reconstruct polynomials exactly, similar to RePU neural networks, and achieves the same universal approximation property for the class of differentiable functions as shown in [15]. Additionally, we experimentally verify that, beyond this theoretical foundation, neural networks with the MRePU activation function possess an appropriate inductive bias when approximating certain classes of functions.
5.1 Approximation Theorems for MRePU
In [15], it is shown that a RePU network of $p$-th order can accurately represent an expanded univariate polynomial using Horner’s method, utilizing $4p$ nodes for multiplication and $2p$ nodes to construct a $p$-th order polynomial. Furthermore, by applying these univariate polynomial results inductively, the representation of multivariate polynomials can also be achieved. Therefore, for the MRePU of order $p$ network, if we can demonstrate that it can construct any $p$-th order polynomial using $2p$ nodes, as in the case of RePU, we obtain the same result. The following lemma formalizes this.
Lemma 7 A polynomial of degree $p$, $\sum_{k=0}^{p}a_{k}x^{k}$ can be approximated by $p$-th order MRePU network with one hidden layer and nodes no more than $2p$.
Proof Consider the neural network constructed as follows.
$y=W_{1}^{T}\sigma_{m;p}(W_{0}^{T}x+b_{0})+v_{0}$
Here, each parameter is defined as follows:
$W_{0}=\begin{bmatrix}1\\ -1\\ \vdots\\ 1\\ -1\end{bmatrix}\in\mathbb{R}^{2p},\quad b_{0}=\begin{bmatrix}u_{1}-1\\ -u_{1}-1\\ \vdots\\ u_{p}-1\\ -u_{p}-1\end{bmatrix}\in\mathbb{R}^{2p},\quad W_{1}=\begin{bmatrix}v_{1}\\ (-1)^{p}v_{1}\\ \vdots\\ v_{p}\\ (-1)^{p}v_{p}\end{bmatrix}\in\mathbb{R}^{2p}.$
In this case, since the components of $W_{0}^{T}x+b_{0}$ where the MRePU activation is applied are symmetric with respect to $x=-1$, only one set of components—either the even-indexed or odd-indexed—remains active. If the odd components remain active, the final output takes the following form:
$v_{k}(x+u_{k})^{p}(x+u_{k}-1).$
In the case of the even components, the output is as follows:
$v_{k}(x+u_{k})^{p}(-x-u_{k}-1).$
In both cases, after binomial expansion, the expression can be written as follows, where $s$ is a value determined by the parity of the index:
$v_{k}\Big{(}(-1)^{s}\sum_{j=0}^{p+1}\binom{p+1}{j}u_{k}^{p+1-j}x^{j}-\sum_{j=0}^{p}\binom{p}{j}u_{k}^{p-j}x^{j}\Big{)}.$
Then, by comparing the final expression of the neural network with the coefficients of the polynomial we aim to construct, $\sum_{k=0}^{p}a_{k}x^{k}$, for $k<p+1$, we obtain the following result:
$\Big{(}(-1)^{s}\binom{p+1}{k}-\binom{p}{k}\Big{)}\sum_{l=1}^{p}(u_{l}^{p+1-k}-u_{l}^{p-k})v_{l}=a_{k}.$
This can be expressed as a matrix equation as follows:
$\begin{bmatrix}u_{1}-1&\dots&u_{p}-1&0\\ \vdots&&\vdots&\vdots\\ u_{1}^{p}-u_{1}^{p-1}&\dots&u_{p}^{p}-u_{p}^{p-1}&0\\ u_{1}^{p+1}-u_{1}^{p}&\dots&u_{p}^{p+1}-u_{p}^{p-1}&1\par\end{bmatrix}\begin{bmatrix}v_{1}\\ \vdots\\ v_{p}\\ v_{0}\end{bmatrix}=\begin{bmatrix}a_{p}\Big{(}(-1)^{s_{p}}\binom{p+1}{p}-\binom{p}{p}\Big{)}^{-1}\\ \vdots\\ a_{1}\Big{(}(-1)^{s_{1}}\binom{p+1}{1}-\binom{p}{1}\Big{)}^{-1}\\ a_{0}\Big{(}(-1)^{s_{0}}\binom{p+1}{0}-\binom{p}{0}\Big{)}^{-1}\end{bmatrix}.$ (9)
Then, the determinant of the top-left $p\times p$ sub-matrix of this equation is given by:
$\prod_{k=1}^{p}(u_{k}-1)\det\bigg{(}\begin{bmatrix}1&\dots&1\\ u_{1}&&u_{p}\\ \vdots&&\vdots\\ u_{1}^{p-1}&\dots&u_{p}^{p-1}\end{bmatrix}\bigg{)}$
The arguments of the determinant on the right form a Vandermonde matrix, so equation (9) has a solution if each $u_{i}$ is distinct and not equal to 1. To make the $p+1$-th order term vanish, it suffices to show the coefficient $(-1)^{s_{1}}v_{1}+\dots+(-1)^{s_{p}}v_{p}$ can be reduced. It can be easily verified that the hyperdeterminant of the Jacobian of the top-left $k\times k$ sub-matrix of equation (9) is a constant multiple of the determinant of the Vandermonde matrix. Therefore, by the inverse function theorem, we can conclude that the following function occupies a positive measure in a neighborhood containing the origin in $\mathbb{R}^{p}$ space.
$F(u_{1},\dots,u_{p})=\begin{bmatrix}u_{1}-1&\dots&u_{p}-1\\ \vdots&&\vdots\\ u_{1}^{p}-u_{1}^{p-1}&\dots&u_{p}^{p}-u_{p}^{p-1}\par\end{bmatrix}^{-1}\begin{bmatrix}a_{p}\Big{(}(-1)^{s_{p}}\binom{p+1}{p}-\binom{p}{p}\Big{)}^{-1}\\ \vdots\\ a_{1}\Big{(}(-1)^{s_{1}}\binom{p+1}{1}-\binom{p}{1}\Big{)}^{-1}\end{bmatrix}.$
Therefore, by choosing sufficiently small $v_{i}$’s, the coefficient of the $p+1$-th order monomial can be reduced to the desired degree. $\square$
Then, by following this lemma and the proof outlined in [15], we obtain the following theorem:
Theorem 8 if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a polynomial of $d$ variables with degree $N$, then $f$ can be exactly reconstructed via MRePU neural network with $12N-$ 1 hidden layers, $(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d})$ number of neurons, $(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d})$ number of parameters and width $12pN^{d-1}+6p(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1})$.
Additionally, as a corollary to the above theorem, we obtain a universal approximation theorem for the class of differentiable functions. Before explaining this, we define a reference norm that will serve as the standard. Definition 9 We define the norm of $f\in C^{s}(A)$, where $A\subset\mathbb{R}^{d}$, as following:
$\|f\|_{C^{s}}:=\sum_{|\alpha|\leq s}\sup_{x\in A}|D^{\alpha}f(x)|,$
where $\alpha$ is an index vector for partial derivative and $|\alpha|$ is total order of derivative.
Corollary 10 Let $f$ be a function belongs to $C^{s}(K)$ where $K\subset\mathbb{R}^{d}$ is a compact set. For any $N\in\mathbb{N}^{+}$, there exists a MRePU neural network $\psi_{N}$ width depth $\mathcal{D}$, width $\mathcal{W}$, number of neurons $\mathcal{U}$ and total number of parameters $\mathcal{P}$ with following specification:
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=12pN^{d-1}+6p(N^{d-1}-N)/(N-1),\\ &\mathcal{U}=(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1),\\ &\mathcal{P}=(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1).\end{split}$
such that for $|\alpha|\leq\min\{s,N\}$, we have
$sup_{K}|D^{\alpha}(f-\psi_{N})|\leq C_{p,s,d,K}\|f\|_{C^{|\alpha|}}N^{-(s-|\alpha|)},$
where $C_{p,s,d,K}$ is a positive constant which depends only on $p,d,s$ and the diameter of $K$.
5.2 Experiments
In this section, we experimentally confirm that the MRePU network, by restricting the hypothesis class and thus providing inductive bias, achieves faster convergence of training loss compared to networks constructed with conventional activations such as ReLU or GELU. RePU is excluded from the comparison group as stacking layers with it is entirely infeasible.
Regression on polynomials and differentiable functions First, we validate the efficiency of MRePU on polynomial functions. The neural network architecture consists of five hidden layers, each with a width of 32, and the input dimension is set according to the variables in each polynomial. For univariate polynomial problems, the training dataset was constructed with inputs from -2.0 to 2.0, divided into 101 points at intervals of 0.04. For $x^{3}-x$, the learning rate was set to 0.005, and $C_{W}$ was set to 0.5 for MRePU and 1.0 for other cases. For $x^{5}-5x^{3}+6x$, the learning rate was set to 0.0001, with $C_{W}$ set to 0.8 for MRePU case and learning rate to 0.001, $C_{W}$ to 1.0 for the others. For multivariate cases, the dataset for $xy$ was created by forming a meshgrid over the range $[-2,2]\times[-2,2]$, with each axis divided into 101 points. Similarly, for $xyz$, a meshgrid was formed over the range $[-1,1]\times[-1,1]\times[-1,1]$, with each axis divided into 51 points. Training was conducted up to 10,000 epochs for all cases. The learning rate for multivariate cases was set to 0.001, and $C_{W}$ was set to 0.55 for $xy$, 0.6 for $xyz$, and 1.0 for the other cases. Each hyperparameter was empirically determined to achieve optimal performance.
$x^{3}-x$ $x^{5}-5x^{3}+6x$ $xy$ $xyz$
MRePU $\bf{4.0*10^{-4}}$ $\bf{2.0*10^{-4}}$ $\bf{2.7*10^{-5}}$ $\bf{2.5*10^{-4}}$
ReLU $8.0*10^{-4}$ $7.0*10^{-4}$ $4.0*10^{-4}$ $4.2*10^{-2}$
GELU $7.2*10^{-2}$ $2.1*10^{-3}$ $4.0*10^{-4}$ $4.2*10^{-2}$
Table 4: Mean final training losses over 25 trials for polynomial approximation for each of the MRePU (order $p=2$), ReLU, and GELU networks. The values in bold represent the lowest values in each column.
Table 4 presents the average final loss over 25 experiments, where neural networks were trained to regress each polynomial function. As observed in the table, MRePU-based neural networks consistently outperform the others in all cases. Notably, for approximating $xyz$, other networks fail to learn effectively, while only the MRePU network achieves meaningful learning. The loss plots for $x^{5}-5x^{3}+6x$ and $xy$ are depicted in Figure 9. Figure 10 illustrates the regression results of $x^{5}-5x^{3}+6x$ using ReLU and MRePU. A key observation is that the trained MRePU network approximates not only the function itself but also its derivative, whereas the ReLU network produces a piecewise function with noticeable discontinuities. Similarly, Figure 11 shows the contour graphs of trained networks with different activation functions for $xy$. The MRePU network demonstrates a highly accurate approximation, while the ReLU network exhibits significant instability in the contours, and GELU shows reduced accuracy.
Figure 9: The training loss plots for each neural network over epochs for the cases of $x^{5}-5x^{3}+6x$ (Left) and $xy$ (Right).
Figure 10: Top: The results of training for regressing $x^{5}-5x^{3}+6x$, where the left plot shows the output of the MRePU with $p=2$ network, and the right plot shows the output of the ReLU network. Bottom: The comparison between the derivative of the trained networks and the actual derivative function, $65x^{4}-15x^{2}+$ 6.
Figure 11: Contour plots for the reference function and the trained neural networks. Top left: MRePU with $p=2$ network. Top right: Reference ($xy$). Bottom left: GELU network. Bottom right: ReLU network.
Next, we conducted experiments on differentiable functions. The neural network architecture remained identical to that used for polynomial regression. The experiments involved regression on $\lfloor x\rfloor_{+}^{3}$ and the functions $f(x)$ and $g(x,y)$, defined as follows.
$f(x)=\begin{cases}x^{4}-3x^{2}+1&\text{if }x\leq-1\\ -x^{2}&\text{if }-1\leq x\leq 0\\ x^{3}&\text{if }0\leq x\leq 1\\ 3x-2&\text{if }1\leq x\end{cases}$
$\begin{split}g(x,y)&=\begin{cases}10(r^{2}-\frac{1}{2})^{2}&\text{if }r\leq\frac{1}{\sqrt{2}}\\ (r^{2}-\frac{1}{2})^{3}&\text{otherwise}\end{cases}\\ r&=\sqrt{x^{2}+y^{2}}\end{split}$
The training dataset for univariate functions was constructed in the same manner as for the polynomials. For multivariate functions, a meshgrid was generated by dividing the domain $[-1,1]\times[-1,1]$ into 51 intervals along each axis. The learning rate and $C_{W}$ values for these experiments were as follows: For $\lfloor x\rfloor_{+}^{3}$, the MRePU network used a learning rate of 0.0004 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $f(x)$, The MRePU network employed a learning rate of 0.001 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $g(x,y)$, the MRePU network used a learning rate of 0.002 and $C_{W}=0.8$, whereas the remaining networks used 0.003 and $C_{W}=1.0$. Each hyperparameter was empirically determined to achieve optimal performance.
$\lfloor x\rfloor_{+}^{3}$ $f(x)$ $g(x,y)$
MRePU $8.1*10^{-5}$ $7.9*10^{-5}$ $4.0*10^{-4}$
ReLU $\bf{3.2*10^{-5}}$ $\bf{5.6*10^{-5}}$ $\bf{2.0*10^{-4}}$
GELU $2.4*10^{-3}$ $8.0*10^{-4}$ $4.4*10^{-1}$
Table 5: Mean final training losses over 25 trials for the approximation of differentiable functions for each of the MRePU (order $p=2$), ReLU, and GELU networks. The values in bold represent the lowest values in each column.
As shown in Table 5 and Figure 12, for differentiable functions with discontinuities in higher-order derivatives, the ReLU network generally demonstrates the best performance. However, as observed in Figures 13 and 14, even though the training loss of the ReLU network is lower (i.e., its $L^{2}$ norm is smaller), the MRePU network exhibits more stable function approximation in terms of contour smoothness and the behavior of the differentiated function. For the GELU network, it is evident that learning fails entirely for $g(x,y)$.
Figure 12: The training loss plots for each neural network over epochs for the cases of $f(x)$ (Left) and $g(x,y)$ (Right).
Figure 13: Top: The results of training for regressing $f(x)$, where the left plot shows the output of the MRePU with $p=2$ network, and the right plot shows the output of the ReLU network. Bottom: The comparison between the derivative of the trained networks and the actual derivative function.
Figure 14: Contour plots for the reference function and the trained neural networks. Top left: MRePU with $p=2$ network. Top right: Reference ($g(x,y)$). Bottom left: GELU network. Bottom right: ReLU network.
Case of Physics-Informed Neural Networks (PINNs) Next, we compare and contrast neural networks constructed with different activation functions using the physics-informed neural networks (PINNs) [26] technique, which solves PDE problems by incorporating the PDE itself into the loss function. The PDE used in this experiment is the Burgers equation, given as:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}}$
We defined the problem domain as $(x,t)\in[-1,1]\times[0,1]$ with $\nu=\frac{0.01}{\pi}$. The initial condition (the function value at $t=0$) was set to $-\sin{\pi x}$, and the boundary condition was defined as $u(-1,t)=u(1,t)=0$. The PINN loss comprises the $L^{2}$ loss for the boundary condition and the loss for the PDE. The number of points used for loss calculations was as follows: 10,000 randomly sampled collocation points within the domain and 100 points for the boundary condition. The hyperparameter settings included the use of the Adam optimizer with a learning rate of 0.001, betas set to $(0.9,0.999)$, and weight decay set to 0. For $C_{W}$, the value was set to 0.45 for MRePU with $p=2$ and 1.0 for all other cases. The neural network architecture was consistent across experiments, with an input dimension of 2, seven hidden layers, and a width of 64 per layer.
Figure 15: The PINN training results of the MRePU with $p=2$ network. Top: The learned neural network function represented as a heatmap over the entire domain. Bottom: Plots of the function cross-sections at $t=0.25$, $t=0.50$, and $t=0.75$.
Figure 16: The PINN training results of the ReLU network. Top: The learned neural network function represented as a heatmap over the entire domain. Bottom: Plots of the function cross-sections at $t=0.25$, $t=0.50$, and $t=0.75$.
The results of the experiment are presented in Figure 15, 16, and 17. Interestingly, unlike the previous cases with polynomials and differentiable functions, the GELU activation exhibited the best performance, followed by the MRePU network. The ReLU network, however, failed to achieve meaningful learning. Additional experiments with the Tanh activation showed a performance similar to that of the GELU network. From this experiment, it can be concluded that the MRePU network is capable of achieving meaningful learning even under the highly complex training dynamics of the PINN loss [27], [28] and with a deep architecture consisting of 7 hidden layers (9 layers in total).
Figure 17: The training loss plots for each neural network over epochs for the case of Burgers equation PINN problem.
To conclude the results of the experiments, it was observed that the characteristics of each activation function (piecewise linear, differentiable, smooth) introduced an inductive bias to the neural network architecture, affecting its effectiveness in approximating specific function classes. These classes included polynomial functions, differentiable functions, and solutions to PINNs (generally smooth functions). Depending on the nature of the task, certain activation functions and architectures were found to be more effective, while others were not. The effectiveness of each activation function, based on the training loss, is summarized as follows:
Polynomial: MRePU $>$ ReLU $>$ GELU
Differentiable function: ReLU $>$ MRePU∗∗ $>$ GELU
PINN for Burgers equation: GELU $>$ MRePU $>$ ReLU
Here, the symbol indicates cases where the training either fails entirely or is effective only under specific conditions. Additionally, the ∗∗ symbol denotes the best effectiveness when considering derivatives as well.
6 Conclusion
In this paper, we analyzed neural networks using Effective Field Theory to theoretically predict and experimentally verify why neural network architectures with RePU activation fail when the network depth increases. Furthermore, we proposed the MRePU activation function, which improves upon RePU, and estimated its effectiveness theoretically. We then empirically confirmed its feasibility by comparing the experimental results with the theoretical predictions. The proposed MRePU was able to facilitate learning even in deep networks without adopting structures like ResNet. Mathematically, we confirmed that MRePU retains the favorable properties of RePU, such as p-differentiability and universal approximation properties. This suggests that the beneficial characteristics proposed in previous studies on RePU can also be extended to MRePU. Moreover, when attempting to approximate a function with a specific differentiability class, MRePU can be used to introduce inductive bias into deep neural networks, thereby enabling effective learning.